### v1

@@ 2.3 Executable outcomes in reaction-function games
    Any outcome is preferred to the case in which there is no fixed point.
@@
The first assumption captures player optimism, familiar for instance from games with externalities (e.g. Aumann and Peleg, 1960; Shenoy, 1979). The second assumption has to do with that fixed points need not exist: one can carefully construct profiles in which players “miscoordinate” to eliminate fixed points (e.g., best-reply reactions in the game of Matching pennies, see Table 3 later).
@@
To summarize, a reaction-function game $(N,\mathcal{R},(U_{i})_{i})$ comprises players $N$, each player $i$ with reaction functions $\mathcal{R}_{i}$ and payoff function $U_{i}\colon\mathcal{R}\to\mathbb{R}\cup\{-\infty\}$ induced by payoffs $u_{i}$ on $A$ and determined by $i$’s preferred fixed point when such a point exists.
@@ 2.4 Reaction-function equilibrium
As players evaluate profiles by their subjectively preferred fixed point, different players may evaluate the same profile by different outcomes. To alleviate any concerns to this point, we will focus on unambiguous profiles, for which, constrained to the set of fixed points, there is a Pareto-superior one.
@@
Table 1: Payoff matrix for Battle of the sexes. The “match other” reactions, which coincide with best replies, are highlighted. The two fixed points are not Pareto ordered, so the profile is not unambiguous.
We take inspiration from Nash equilibrium as defined in Subsection 2.1 to formulate a generalized solution concept for reaction-function games. In equilibrium, no unilateral reaction-function change is beneficial. The restriction to unambiguous profiles aligns with the Nash equilibrium requirement that players’ beliefs about the consequences of their actions are correct: all players associate to an equilibrium profile $R$ the same outcome $a\in\mathcal{E}(R)$.
@@
If players are restricted to constant reaction functions, then fixed points are unique (Remark 1) and reaction-function equilibrium is logically equivalent to Nash equilibrium. We record this in Observation 1. Applied to the Battle of the sexes, there are two such constant reaction-function equilibria, namely that both players always play $x$, or both always play $y$ [^3] For the Battle of the sexes, we find the striking conclusion that any outcome can be supported in reaction-function equilibrium, including miscoordination; see Theorem 3.2. Whereas the profile of best replies, BR, is not unambiguous and thus not a reaction-function equilibrium, the worst replies actually support miscoordination, that is $(x,y)$ or $(y,x)$. This is an example of a “bad equilibrium” in which each player rationalizes playing badly by that the other does so as well. This issue is inherited directly from Nash equilibrium (e.g. the Pareto-inferior Nash equilibrium in the Divide the dollar-game, see Brams and Taylor, 1994) but gets amplified as reaction-function games are more complex.
@@
To conclude that a profile is not a reaction-function equilibrium, it suffices to consider constant deviations: if a player can beneficially deviate through some reaction function, then there is also a constant reaction that the player can deviate through. If player $i$ targets the preferred fixed point $a^{\prime}\in\mathcal{E}(R^{\prime}_{i},R_{-i})$ by deviating through $R^{\prime}_{i}$, then $i$ also beneficially deviates by always playing $a^{\prime}_{i}$ (as then $a^{\prime}$ again is a fixed point). This parallels mixed-strategy equilibrium, where it analogously suffices to consider pure-strategy deviations. Example 1 illustrates this point and shows that the best-reply profile need not be a reaction-function equilibrium and that always playing a dominated action can be superior to playing the dominating action.
@@
Consider the two-player game with payoffs in Table 2. Best replies are highlighted: we have $\textit{BR}_{1}(x)=a$ and $\textit{BR}_{1}(y)=b$ whereas action $x$ is dominant for the second player, $\textit{BR}_{2}(a)=\textit{BR}_{2}(b)=x$.
@@
There is a unique fixed point, $\mathcal{E}(\textit{BR})=\{(a,x)\}$. However, BR is not a reaction-function equilibrium as player $2$ prefers to deviate through the constant reaction $R^{\prime}_{2}$ pertaining to the dominated action $y$, $R^{\prime}_{2}(a)=R^{\prime}_{2}(b)=y$: we have $\mathcal{E}(\textit{BR}_{1},R^{\prime}_{2})=\{(b,y)\}$, so $U_{2}(\textit{BR}_{1},R^{\prime}_{2})=2>1=U_{2}(\textit{BR})$. Indeed, $(\textit{BR}_{1},R^{\prime}_{2})$ is a reaction-function equilibrium. For contrast, by Observation 1, it suffices that both play constant reactions to support the Nash equilibrium $(a,x)$ in reaction-function equilibrium. $\circ$
There are often many reaction-function equilibria supporting the same outcome. This mirrors how a wide range of off-equilibrium threats and promises can support the same equilibrium path in infinitely repeated games. It will at times be more interesting to explore which outcomes can be supported rather than the particular reaction functions used to support them. Formally, outcome $a$ is supported in reaction-function equilibrium if there exists a reaction-function equilibrium $R\in\mathcal{R}$ such that $a\in\mathcal{E}(R)$ and $u_{i}(a)\geq u_{i}(a^{\prime})$ for each player $i$ and outcome $a^{\prime}\in\mathcal{E}(R)$.
@@ 3.1 Equilibrium existence
Our first result records that reaction-function equilibria always exist. This is in contrast to Nash equilibrium in the corresponding strategic game (recall that players cannot mix). By Observation 1, this implies that we will have to go beyond constant reaction functions to prove reaction-function equilibrium existence. We formalize the argument below, which is instructive and of independent interest as it links to yet another class of games, namely sequential-move games. These neatly bridge simultaneous-move and reaction-function games.
Suppose players take action in sequence $1,\dots,n$, each observing all prior choices before making theirs (in the conventional sense, not using reaction functions). This defines a finite extensive-form game of perfect information. Such games do have Nash equilibria (Kuhn, 1953, use backward induction). This equilibrium is a strategy profile $s=(s_{1},\dots,s_{n})$ where each $s_{i}\colon A_{<i}\to A_{i}$ maps an action $s_{i}(a_{1},\dots,a_{i-1})\in A_{i}$ to all prior moves $(a_{1},\dots,a_{i-1})\in A_{<i}$. This translates into the reaction function $R_{i}\in\mathcal{R}_{i}$ with $R_{i}(a_{-i})=s_{i}(a_{<i})$ for all $a_{-i}\in A_{-i}$. As sketched in Remark 1, there is a unique fixed point at $R$, namely the outcome realized along the equilibrium path of the extensive-form game, and the same holds at $(R^{\prime}_{i},R_{-i})$ for all constant reactions $R^{\prime}_{i}\in\mathcal{R}_{i}$. If $i$ were able to deviate in the reaction-function game by always taking action $a^{\prime}_{i}$, then $i$ would also deviate in the extensive-form game through strategy $s^{\prime}_{i}$ with $s^{\prime}_{i}(a_{<i})=a^{\prime}_{i}$ for each $a_{<i}\in A_{<i}$. Therefore, outcomes supported in Nash equilibrium in the extensive-form game—which exist—are also supported in reaction-function equilibrium. [^4] One may wonder whether the converse is true, that is, if outcome $a$ is supported in reaction-function equilibrium, can we order the players such that $a$ is realized on the equilibrium path of the corresponding extensive-form game? This is not true. The “bad equilibria” of the Battle of the sexes (see footnote 3) cannot be supported in this way as they require a “cyclical” reaction-function profile (compare Remark 1).
@@
If we apply this construction to Matching pennies (Table 3), we find profiles that comprise exactly one constant and one best-reply reaction (e.g., $R_{1}(H)=R_{1}(T)=H$ and $R_{2}=\textit{BR}_{2}$ with $R_{2}(H)=T$ and $R_{2}(T)=H$). These are indeed the only reaction-function equilibria.
@@ 3.2 Maxmin payoff bound for two-player games
Our next result mimics the “folk theorems” obtained in Tennenholtz (2004) and Kalai et al. (2010) but restricted to two-player games. Theorem 3.2 asserts that an outcome can be supported precisely when it awards every player at least their maxmin payoff (or “security level”, Maschler et al., 2013). Player $i$’s maxmin payoff $\underaccent{\bar}{v}_{i}\in\mathbb{R}$ is the highest payoff $i$ can ensure irrespective of the actions of the others:
$\underaccent{\bar}{v}_{i}=\max_{a_{i}\in A_{i}}\min_{a_{-i}\in A_{-i}}u_{i}(a_{i},a_{-i}).$
@@
Theorem 3.2 parallels folk theorems of infinitely repeated games but is now obtained in a one-shot setting. The “trigger strategy” of repeated games here maps into a “promise and threat” reaction function. To support outcome $a^{*}\in A$ (with at least maxmin payoffs), play $R_{i}(a^{*}_{-i})=a^{*}_{i}$ and, for $a_{-i}\neq a^{*}_{-i}$, select $a_{i}=R_{i}(a_{-i})$ to minimize the other player’s payoff $u_{-i}(a)$. Compared to the existence proof of Theorem 3.1, which involved at least one constant reaction, we now go yet further and all players choose more complex reactions. The proof is in the Appendix.
@@
To illustrate Theorem 3.2, consider the Prisoner’s dilemma (Table 4). In the strategic game, the Pareto-efficient cooperative outcome $(C,C)$ cannot be sustained as “defecting” ($D$) is dominant for each player. However, the cooperative outcome can be supported in the reaction-function game, and in precisely one way: each player cooperates if and only if the other player does (highlighted in Table 4). Such reactions lead to two Pareto-ordered fixed points, the inferior Nash equilibrium $(D,D)$ and the superior outcome $(C,C)$. Hence, compared to the unique fixed point in the construction offered in Theorem 3.1, we may now obtain reaction-function equilibria with multiple (but still Pareto-ordered) fixed points.
@@ 3.3 More players, more possibilities for support
Once there are more players, we can carefully construct reactions to support an outcome by out-ruling any other outcome from ever becoming a fixed point even following unilateral deviations. That is, for each outcome $a$, we can find a profile $R\in\mathcal{R}$ such that, for each deviation $R^{\prime}_{i}\in\mathcal{R}_{i}$, we have $\mathcal{E}(R^{\prime}_{i},R_{-i})\subseteq\mathcal{E}(R)=\{a\}$ and thus $U_{i}(R^{\prime}_{i},R_{-i})\leq U_{i}(R)$. This illustrates how an outcome can be supported simply because there is no other alternative. The proof is a combinatorial exercise found in the Appendix. For the case with three players and two actions, we can find examples in which some outcomes cannot be supported in reaction-function equilibrium (see again the Appendix).
@@
The construction employed in the proof of Theorem 3.3 is highly artificial and provides little in terms of practical insight—for instance, player reactions are set completely independent from their payoffs. In what follows, we refine the solution concept to simultaneously make equilibria more practically compelling as well as eliminate support for “bad” outcomes.
@@
When constructing reaction functions, some reactions appear more unsafe than others. For instance, Theorem 3.2 builds on “promise and threat” reactions, where an outcome with at least maxmin payoffs is supported on the premise of maximal punishment for any deviation. Even though these “threats” are not carried out in equilibrium, it seems risky for player $i$ to base their reaction solely on player $j$’s payoffs. Next, we address this by introducing the idea of safe play.
“Safe play” in strategic games entails taking a safe action $\underaccent{\bar}{a}_{i}$ to ensure at least maxmin payoff $\underaccent{\bar}{v}_{i}$ regardless the actions of the others (e.g. Maschler et al., 2013). This individual risk-minimizing behavior can lead to collectively bad outcomes, for instance, in Prisoner’s dilemma (Table 4). With reaction functions, there are more options. We say that reaction $R_{i}$ is safe if, for all actions $a_{-i}$ of the others, the outcome $(R_{i}(a_{-i}),a_{-i})\in A$ awards $i$ at least payoff $\underaccent{\bar}{v}_{i}$. [^5] A stronger requirement would be $U_{i}(R_{i},R_{-i})\geq\underaccent{\bar}{v}_{i}$ for each $R_{-i}\in\mathcal{R}_{-i}$, but this is too demanding. With more than two players, there is $R_{-i}\in\mathcal{R}_{-i}$ such that $\mathcal{E}(R)=\emptyset$ and thus $U_{i}(R)=-\infty$ for all $R_{i}\in\mathcal{R}_{i}$. A reaction-function equilibrium $R$ is safe if each player’s reaction $R_{i}$ is safe.
@@
Safe reaction-function play is always possible, for instance by playing best replies $\textit{BR}_{i}$. Yet simpler is the constant reaction function to always play a safe action $\underaccent{\bar}{a}_{i}$. If the resulting action profile $\underaccent{\bar}{a}$ of such play is a Nash equilibrium in the strategic game, then the corresponding constant reactions form a reaction-function equilibrium (Observation 1). In cases like a “race to the bottom”, where players say won’t exert effort unless others do, the low-effort outcome may be a Nash equilibrium and thus supported in safe reaction-function equilibrium. But importantly, if Pareto improvements are possible, they too can be supported in safe reaction-function equilibrium. Hence, whenever a safe reaction-function equilibrium exists, a Pareto-efficient one does as well.
@@
If a safe outcome $\underaccent{\bar}{a}$ of the strategic game is a Nash equilibrium, then there is a safe reaction-function equilibrium. In general, if a safe reaction-function equilibrium exists, then there exists a Pareto-efficient safe reaction-function equilibrium.
To illustrate Theorem 4.1 and how reaction functions make a difference compared to strategic games, consider the Stag hunt (tracing back to Rousseau’s A Discourse on Inequality) in Table 5. In the strategic game, both taking safe action $y$ leads to the risk-dominant Nash equilibrium (Harsanyi and Selten, 1988). In contrast, matching the action of the other as highlighted is safe and supports the superior payoff-dominant outcome $(x,x)$.
@@
Thus, safe reaction-function equilibrium ensures a high-enough lower bound on player payoffs without having to compromise in achieving yet better outcomes. However, existence is not guaranteed. Example 1 already showed that all playing safe need not be an equilibrium, and Example 2 goes further to show that there need not exist a safe equilibrium.
@@
Example 2 gives further insights on our structural results in Section 3. In consequence, the construction for our existence result (Theorem 3.1) involves unsafe play and our two-player characterization (Theorem 3.2) does not extend to safe reactions. It is by definition that “bad” outcomes awarding less than maxmin payoffs cannot be supported in safe play; hence, safe play addresses at least the most problematic aspect of equilibrium multiplicity (Theorem 3.3). Most reaction functions that we explore next, in Section 5, will be safe.

### v2

@@ 2.3 Fixed points in reaction-function games
    All fixed-point outcomes are preferred to the case in which there is no fixed point.
@@
The first assumption captures player optimism, familiar for instance from games with externalities (e.g. Aumann and Peleg, 1960; Shenoy, 1979). The second addresses profiles $R$ for which $\mathcal{E}(R)=\emptyset$. Such “no fixed point” profiles can arise (e.g., best-reply reactions in the game of Matching pennies, see Table 3 later) and we need a convention for how players evaluate them. Our choice $U_{i}(R)=$-$\infty$ should be understood purely as a normalization: it means that a profile without fixed points cannot be implemented and is therefore ranked below profiles with fixed points. Importantly, -$\infty$ does not represent an actual “punishment” imposed by any player or by the mechanism; it is simply a convenient way to encode that, absent a fixed point, the model does not specify a realized outcome. In some contexts, this ranking is perfectly suitable: for instance, in a surplus-sharing or bargaining setting, any implementable agreement may be preferred by all to failing to agree. The assumption is not needed to establish equilibrium existence (Theorem 3.1) nor is it necessary if one instead imposes suitable restrictions on the permissible reaction functions (compare Section 5).
@@
To summarize, a reaction-function game $(N,\mathcal{R},(U_{i})_{i})$ comprises a set of players $N$. Each player $i$ has access to a set of reaction functions $\mathcal{R}_{i}$ and is endowed with a payoff function $U_{i}\colon\mathcal{R}\to\mathbb{R}\cup\{-\infty\}$. This is induced by the payoffs $u_{i}$ on $A$ and specifically determined by $i$’s preferred fixed point when such a point exists.
@@ 2.4 Reaction-function equilibrium
Next, we develop our solution concept for reaction-function games. As players evaluate profiles by their subjectively preferred fixed point, different players may evaluate the same profile by different outcomes. We will rule out this in equilibrium. That is, an equilibrium profile $R$ will be unambiguous, meaning that there is a Pareto-superior fixed point in $\mathcal{E}(R)$ and hence a clear common outcome to associate to the profile. This is not a restriction on what players can play but rather a strengthening of the solution concept to be defined next (akin to restricting to symmetric equilibria in symmetric games). [^2] There are other ways to achieve the same effect: for instance, one could assume players have agreed beforehand on a “selection function” that mechanically selects a fixed point from $\mathcal{E}(R)$ or explicitly model player beliefs about what outcome will be realized. Our approach is simple, interpretable, and quite restrictive; the other approaches could lead to yet more equilibria. As we shall see, equilibrium multiplicity is already a concern with our stronger assumption.
@@
Table 1: Payoff matrix for Battle of the sexes. The “match other” reactions, which coincide with best replies, are highlighted. The two fixed points are not Pareto ordered, so the profile is not unambiguous and, in consequence, not a reaction-function equilibrium.
We take inspiration from Nash equilibrium as defined in Subsection 2.1 to formulate a generalized solution concept for reaction-function games. In such an equilibrium, no unilateral reaction-function change is beneficial. The restriction to unambiguous profiles aligns with the Nash equilibrium requirement that players’ beliefs about the consequences of their actions are correct: all players associate to an equilibrium profile $R$ the same outcome $a\in\mathcal{E}(R)$.
@@
If players are restricted to constant reaction functions, then fixed points are unique for each profile (Remark 1) and reaction-function equilibrium is logically equivalent to Nash equilibrium. We record this in Observation 1. Applied to the Battle of the sexes, there are two such constant reaction-function equilibria: both players always play $x$, or both always play yy. [^3] For the Battle of the sexes, we find the striking conclusion that any outcome can be supported in reaction-function equilibrium, including miscoordination; see Theorem 3.2. Whereas the profile of best replies, BR, is not unambiguous and thus not a reaction-function equilibrium, the worst replies actually support miscoordination, that is $(x,y)$ or $(y,x)$. This is an example of a “bad equilibrium” in which each player rationalizes playing badly by that the other does so as well. This issue is inherited directly from Nash equilibrium (e.g. the Pareto-inferior Nash equilibrium in the Divide the dollar-game, see Brams and Taylor, 1994) but gets amplified as reaction-function games are more complex.
@@
To conclude whether a profile is a reaction-function equilibrium or not, it suffices to consider constant deviations. This is recorded in Observation 2. Indeed, if player $i$ targets the preferred fixed point $a^{\prime}\in\mathcal{E}(R^{\prime}_{i},R_{-i})$ by deviating through $R^{\prime}_{i}$, then $i$ also beneficially deviates through the constant reaction $R^{\prime\prime}_{i}$ comprised of always playing $a^{\prime}_{i}$ as then $a^{\prime}\in\mathcal{E}(R^{\prime\prime}_{i},R_{-i})$. [^4] This parallels mixed-strategy equilibrium, where it analogously suffices to consider pure-strategy deviations.
Observation 2.
A profile $R\in{\cal R}$ is a reaction-function equilibrium if and only if there does not exist a player $i$ and constant reaction function $R^{\prime}_{i}\in\mathcal{R}_{i}$ such that $U_{i}(R^{\prime}_{i},R_{-i})>U_{i}(R)$.
Best-reply reactions are a natural candidate for equilibrium play. However, in reaction-function games, the equilibrium condition is evaluated in terms of the fixed points induced by the profile, and a unilateral deviation can completely change this set. Hence, even if, for instance, the profile of best replies has a unique fixed point, it need not be stable: a player may prefer to deviate (through a constant reaction function)—possibly even corresponding to an action dominated in the strategic game. Example 1 illustrates this mechanism.
@@
Consider the two-player game with payoffs in Table 2. Best replies are highlighted: we have $\textit{BR}_{1}(x)=a$ and $\textit{BR}_{1}(y)=b$ whereas action $x$ dominates $y$ for the second player, $\textit{BR}_{2}(a)=\textit{BR}_{2}(b)=x$.
@@
There is a unique fixed point, $\mathcal{E}(\textit{BR})=\{(a,x)\}$. However, BR is not a reaction-function equilibrium as player $2$ prefers to deviate through the constant reaction $R^{\prime}_{2}$ with $R^{\prime}_{2}(a)=R^{\prime}_{2}(b)=y$: we have $\mathcal{E}(\textit{BR}_{1},R^{\prime}_{2})=\{(b,y)\}$, so $U_{2}(\textit{BR}_{1},R^{\prime}_{2})=2>1=U_{2}(\textit{BR})$. Indeed, $(\textit{BR}_{1},R^{\prime}_{2})$ is a reaction-function equilibrium. $\circ$
There are often many reaction-function equilibria supporting the same outcome. This mirrors how a wide range of off-equilibrium threats and promises can support the same equilibrium path in infinitely repeated games. It will at times be more interesting to explore which outcomes can be supported rather than the particular reaction functions used to support them. Formally, outcome $a$ is supported in reaction-function equilibrium if $a$ is the outcome unambiguously associated to an equilibrium profile; that is, there is a reaction-function equilibrium $R\in\mathcal{R}$ such that $a\in\mathcal{E}(R)$ and $u_{i}(a)\geq u_{i}(a^{\prime})$ for each player $i$ and outcome $a^{\prime}\in\mathcal{E}(R)$.
@@ 3.1 Equilibrium existence
Our next result establishes that reaction-function equilibria always exist. This is in contrast to Nash equilibrium in the corresponding strategic game (recall that players cannot mix). By Observation 1, this implies that we will have to go beyond constant reaction functions to prove reaction-function equilibrium existence. We formalize the argument below, which is instructive and of independent interest as it links to yet another class of games, namely sequential-move games. These neatly bridge simultaneous-move strategic games and reaction-function games.
Suppose players take action in sequence $1,\dots,n$, each observing all prior choices $a_{<i}\in A_{<i}$ before making theirs (in the conventional sense, i.e., not using reaction functions). This defines a finite extensive-form game of perfect information. Such games do have Nash equilibria (Kuhn, 1953, use backward induction). Such an equilibrium is a strategy profile $s=(s_{1},\dots,s_{n})$ where each $s_{i}\colon A_{<i}\to A_{i}$ maps histories of prior moves $(a_{1},\dots,a_{i-1})\in A_{<i}$ to an action $s_{i}(a_{1},\dots,a_{i-1})\in A_{i}$ by player $i$. This translates into the reaction function $R_{i}\in\mathcal{R}_{i}$ with $R_{i}(a_{-i})=s_{i}(a_{<i})$ for all $a_{-i}\in A_{-i}$. As sketched in Remark 1, there is a unique fixed point at $R$, namely the outcome realized along the equilibrium path of the extensive-form game, and the same holds at $(R^{\prime}_{i},R_{-i})$ for all constant reactions $R^{\prime}_{i}\in\mathcal{R}_{i}$. If $i$ were able to deviate in the reaction-function game by always taking action $a^{\prime}_{i}$, then $i$ would also deviate in the extensive-form game through strategy $s^{\prime}_{i}$ with $s^{\prime}_{i}(a_{<i})=a^{\prime}_{i}$ for each $a_{<i}\in A_{<i}$. Therefore, outcomes supported in Nash equilibrium in the extensive-form game—which exist—are also supported in reaction-function equilibrium. [^5] One may wonder whether the converse is true, that is, if outcome $a$ is supported in reaction-function equilibrium, can we order the players such that $a$ is realized on the equilibrium path of the corresponding extensive-form game? This is not true. The “bad equilibria” of the Battle of the sexes (see footnote 3) cannot be supported in this way as they require a “cyclical” reaction-function profile (compare Remark 1).
@@
If we apply this construction to Matching pennies (Table 3), we find profiles that comprise one constant and one best-reply reaction (e.g., $R_{1}(H)=R_{1}(T)=H$ and $R_{2}=\textit{BR}_{2}$ with $R_{2}(H)=T$ and $R_{2}(T)=H$). These are indeed the only reaction-function equilibria.
@@ 3.2 Maxmin payoff bound for two-player games
Our next result mimics the “folk theorems” obtained in Tennenholtz (2004) and Kalai et al. (2010) for two-player games. Theorem 3.2 asserts that an outcome can be supported precisely when it awards every player at least their maxmin payoff (or “security level”, Maschler et al., 2013). Define player $i$’s maxmin payoff $\underaccent{\bar}{v}_{i}\in\mathbb{R}$ as the highest payoff $i$ can ensure irrespective of the actions of the others, i.e.,
$\underaccent{\bar}{v}_{i}\equiv\max_{a_{i}\in A_{i}}\min_{a_{-i}\in A_{-i}}u_{i}(a_{i},a_{-i}).$
@@
Theorem 3.2 parallels folk theorems of infinitely repeated games but is now obtained in a one-shot setting. The “trigger strategy” of repeated games here maps into a “promise and threat” reaction function. To support outcome $a^{*}\in A$ (with at least maxmin payoffs), play $R_{i}(a^{*}_{-i})=a^{*}_{i}$ and, for $a_{-i}\neq a^{*}_{-i}$, select $a_{i}=R_{i}(a_{-i})$ to minimize the other player’s payoff $u_{-i}(a)$. Compared to the existence proof of Theorem 3.1, which involved at least one constant reaction, we now go yet further and both players choose more complex reactions. The proof is in the Appendix.
@@
To illustrate Theorem 3.2, consider the Prisoner’s dilemma (Table 4). In the strategic game, the Pareto-efficient cooperative outcome $(C,C)$ cannot be supported as “defecting” ($D$) is dominant for each player. However, the cooperative outcome can be supported in the reaction-function game, and in precisely one way: each player cooperates if and only if the other player does (highlighted in Table 4). Such reactions lead to two Pareto-ordered fixed points, the inferior Nash equilibrium $(D,D)$ and the superior outcome $(C,C)$. Hence, compared to the unique fixed point in the construction offered in Theorem 3.1, we may now obtain reaction-function equilibria with multiple (but still Pareto-ordered) fixed points.
@@ 3.3 More players, more possibilities for support
Once there are more players, we can carefully construct reactions to support an outcome by ruling out any other outcome from ever becoming a fixed point even following unilateral deviations. That is, for each outcome $a$, we can find a profile $R\in\mathcal{R}$ such that, for each deviation $R^{\prime}_{i}\in\mathcal{R}_{i}$, we have $\mathcal{E}(R^{\prime}_{i},R_{-i})\subseteq\mathcal{E}(R)=\{a\}$ and thus $U_{i}(R^{\prime}_{i},R_{-i})\leq U_{i}(R)$. [^6] Importantly, this uses the assumption $\mathcal{E}(R^{\prime}_{i},R_{-i})=\emptyset\implies U_{i}(R^{\prime}_{i},R_{-i})=$-$\infty$. This illustrates how an outcome can be supported simply because there is no other alternative. The proof is a combinatorial exercise found in the Appendix. For the case with three players and two actions, we can find examples in which some outcomes cannot be supported in reaction-function equilibrium (see again the Appendix).
@@
The construction employed in the proof of Theorem 3.3 is highly artificial and provides little in terms of practical insight—for instance, player reactions are set completely independent from their payoffs. Up to this point, we have imposed little structure on reactions. This helps to establish existence and broad support results, but also allows play that is hard to meaningfully interpret. In what follows, we introduce additional constraints on reaction functions (starting with safe play) to move from theoretical possibility to more compelling predictions and to reduce support for “bad” outcomes.
@@
When constructing reaction functions, some reactions appear less safe than others. For instance, Theorem 3.2 builds on “promise and threat” reactions, where an outcome with at least maxmin payoffs is supported on the premise of maximal punishment for any deviation. Even though these “threats” are not carried out in equilibrium, it seems risky for player $i$ to base their reaction solely on player $j$’s payoffs. Next, we address this by introducing the idea of safe play.
“Safe play” in strategic games entails taking a safe action $\underaccent{\bar}{a}_{i}$ to ensure at least maxmin payoff $\underaccent{\bar}{v}_{i}$ regardless the actions of the others (e.g. Maschler et al., 2013). This individual risk-minimizing behavior can lead to collectively bad outcomes, for instance, in Prisoner’s dilemma (Table 4). With reaction functions, there are more options. We say that reaction $R_{i}$ is safe if, for all actions $a_{-i}$ of the others, the outcome $(R_{i}(a_{-i}),a_{-i})$ awards $i$ at least her maxmin payoff $\underaccent{\bar}{v}_{i}$. [^7] A stronger requirement would be $U_{i}(R_{i},R_{-i})\geq\underaccent{\bar}{v}_{i}$ for each $R_{-i}\in\mathcal{R}_{-i}$, but this is too demanding. With more than two players, there is $R_{-i}\in\mathcal{R}_{-i}$ such that $\mathcal{E}(R)=\emptyset$ and thus $U_{i}(R)=$-$\infty$ for all $R_{i}\in\mathcal{R}_{i}$. A reaction-function equilibrium $R\in{\cal R}$ is safe if each player’s reaction $R_{i}\in{\cal R}_{i}$ is safe.
@@
Safe reaction-function play is always possible, for instance by playing best replies $\textit{BR}_{i}$. Yet simpler is the constant reaction function to always play a safe action $\underaccent{\bar}{a}_{i}$. If the resulting action profile $\underaccent{\bar}{a}$ of such play is a Nash equilibrium in the strategic game, then the corresponding constant reactions form a reaction-function equilibrium (Observation 1). In cases like a “race to the bottom”, where players, say, will not exert effort unless others do, the low-effort outcome may be a Nash equilibrium and thus supported in safe reaction-function equilibrium. But importantly, if Pareto improvements are possible, they too can be supported in safe reaction-function equilibrium. Hence, whenever a safe reaction-function equilibrium exists, a Pareto-efficient one does as well.
@@
If a safe outcome $\underaccent{\bar}{a}\in A$ of the strategic game is a Nash equilibrium, then there is a safe reaction-function equilibrium. In general, if a safe reaction-function equilibrium exists, then there exists a Pareto-efficient safe reaction-function equilibrium.
To illustrate Theorem 4.1 and how reaction functions make a difference compared to strategic games, consider the Stag hunt (tracing back to Rousseau’s A Discourse on Inequality) in Table 5. In the strategic game, both taking safe action $y$ leads to the risk-dominant Nash equilibrium (Harsanyi and Selten, 1988). In contrast, reacting by matching the action of the other (highlighted) is safe and supports the superior payoff-dominant outcome $(x,x)$. Thus, safe reaction-function equilibrium ensures a high-enough lower bound on player payoffs without having to compromise in achieving yet better outcomes.
@@
However, existence is not guaranteed. Example 1 already showed that all playing best replies, i.e., playing safe, need not be an equilibrium. Example 2 goes further to show that there need not exist a safe equilibrium.
@@
Example 2 gives further insights on our main results in Section 3. In consequence, the construction for our existence result (Theorem 3.1) involves unsafe play and our two-player characterization (Theorem 3.2) does not extend to safe reactions. It is by definition that “bad” outcomes awarding less than maxmin payoffs cannot be supported in safe play; hence, safe play addresses at least the most problematic aspect of equilibrium multiplicity (Theorem 3.3). Most reaction functions that we explore next, in Section 5, will be safe.